At financial year end, 31 December 2020, the following appeared in the general ledger of SA Attorneys:


Dr   Insurance expense N1 Cr


 

2020

31/12 Bank CPJ R 41 600

 

 

Additional information:


SA Attorneys entered into an insurance contract with Africa Insurance agency on 1 November 2019. The terms and conditions of the contract stipulates that SA Attorneys should pay an amount of R3 200 on a monthly basis for insurance. The payment is due on the seventh day of each month.

 

What is the amount for the insurance expense that should appear in the trial balance of SA Attorneys for the year ended 31 December 2020?


 A. R3 200

B. R41 600

 C. R38 400

D. R32 000


 

 

 

 

 

 

 

 



1
Expert's answer
2021-06-10T13:14:08-0400

A. R3 200

i.e Accrued expense is given by Total insurance paid - Accumulated insurance expense (R3 200 "\times"12) = R(41600 - 38400) = R3, 200.